OpenStudy (anonymous):

Evaluate f(x) = x² + 1 for f(-1). A.0 B.undefined C.2

OpenStudy (callisto):

C

OpenStudy (anonymous):

c

OpenStudy (callisto):

f(-1) = (-1) ²+ 1 = 1+1 = 2 =C